# Towards sub-milliarcsecond astrometric precision using seeing-limited imaging

### v1

@@ Abstract
The Earth’s atmospheric turbulence degrades the precision of ground-based astrometry. Here we discuss these limitations and propose that, with proper treatment of systematics and by leveraging the many epochs available from the Korean Microlensing Telescope Network (KMTNet), seeing-limited observations can reach sub-milliarcsecond precision. Such observations may be instrumental for the detection of Galactic black holes via microlensing. We present our methodology and pipeline for precise astrometric measurements using seeing-limited observations. The method is a variant of Gaia’s Astrometric Global Iterative Solution (AGIS) that include several detrending steps. Tests on 6,500 images of the same field, obtained by KMTNet with typical seeing condition of 1 arcsecond and pixel scale of 0.4 arcsecond, suggest that we can achieve, at the bright end (mag $\lesssim$17), relative proper motion precision of 0.1-0.2 mas yr-1, over a baseline of approximately five years, using data from the Cerro Tololo Inter-American Observatory (CTIO) site. The precision is estimated using bootstrap simulations and further validated by comparing results from two independent KMTNet telescopes.
@@ 1 Introduction
For the past two millennia, precision astrometry provided a plethora of discoveries. However, due to the Earth’s turbulent atmosphere, ground-based astrometry is considerably less efficient (and even less cost-effective) compared with space-based astrometry. The main problems of ground-based astrometry are: (i) The starting point of the precision (i.e., given a single photon) is the seeing rather than the diffraction limit; (ii) Due to atmospheric scintillations, the photons arriving from the stars are spatially correlated, and therefore the measurement precision decrease slower than $1/\sqrt{N_{\rm ph}}$, where $N_{\rm ph}$ is the number of photons. Although adaptive optics observations and interferometric techniques (e.g., Genzel et al. 1997; Ghez et al. 2005; Gillessen et al. 2010; Dong et al. 2019) can partially solve these problems, they are currently expensive and limited to a small field of view. These two facts make it currently difficult for ground-based astrometry to compete with space-based astrometry. However, even the leading space astrometry mission, Gaia (Gaia Collaboration et al. 2016), has some limitations. Two such limitations are the capability of providing astrometric measurements over short time scales ($\lesssim$ a few months), and the relative inefficiency of Gaia in crowded fields, such as the Galactic bulge. Indeed, several science cases can benefit from precision astrometry measured on time scales shorter than a few months, and/or targets in the Galactic bulge. Examples include searching for astrometric microlensing due to compact objects (e.g., Lu et al. 2016; Sahu et al. 2017; Rybicki et al. 2018; Sahu et al. 2022), searching for lensed quasars and measuring their time delays (e.g., Springer & Ofek 2021a, b), and searching for binary asteroids (e.g., Segev et al. 2023).
@@
In this work, we develop and test a method for precision astrometry using high-cadence, ground-based observations of microlensing events towards the Galactic bulge. Our main motivation for the development of this methodology is to find more examples of isolated stellar-mass black holes (BH), and eventually to measure their mass function. This is of great importance for our understanding of the core-collapse supernova explosion mechanism (e.g., Heger et al. 2003) and the birth of BH binaries and gravitational wave events (e.g., Abbott et al. 2016a). To date, most of the known stellar mass black holes are detected only in binary systems such as X-ray binaries (e.g., McClintock & Remillard 2003 and Özel et al. 2010) and BH-BH (or BH-neutron star) mergers by Gravitational Waves (e.g., Abbott et al. 2016b). Three quiescent stellar-mass black holes were identified through their relatively wide-orbit stellar companions in Gaia observations (El-Badry et al., 2023a, b; Panuzzo et al., 2024). To study the abundance of black holes in the Galaxy using binary systems, a complete understanding of the system’s evolution is required. Unfortunately, the evolution of such binary systems is complicated and likely includes physical processes that are not fully understood. Even with a complete understanding, binary-based methods would remain inherently limited, as they trace only those black holes that remain in observable binaries.
@@
Blaes & Madau (1993) and Agol & Kamionkowski (2002) suggested that isolated compact objects can be detected via X-ray emission that takes place due to Bondi-Hoyle-like accretion from the interstellar medium (ISM). However, searches for such objects have not revealed good candidates so far (Maoz et al. 1997; Chisholm et al. 2003 and Mereghetti et al. 2022). Furthermore, even if such objects are identified, measuring their numbers and masses may prove extremely challenging due to the complexities of accretion physics.
@@
The study of microlensing events is currently the most promising path for detecting isolated stellar-mass black holes. To date, tens of thousands of gravitational microlensing events have been detected by the flux magnification using ground-based telescopes, which observe the Galactic Bulge at high cadence, e.g., OGLE (Udalski et al., 1992); MOA (Abe et al., 2008); KMTNet (Kim et al., 2016). However, the microlensing light curve only rarely provides enough information for the lens mass to be inferred. There are two approaches to deal with this problem. The first is the statistical approach in which, based on the statistical properties of the microlensing population (e.g., the velocity distribution of lenses and sources) one can infer the mass distribution of a sample of microlensing events.(e.g., Gould 2000; Mróz & Wyrzykowski 2021; Mroz et al. 2021; Rybicki et al. 2024). The second approach is breaking the degeneracy and measuring the lens mass. Because, in the simplest case, there are two observables (the magnification and time scale), and five unknowns (lens mass, source distance, lens distance, impact parameter, and relative tangential velocity between the source and the lens), the problem is degenerate. For a Galactic microlensing event, the source distance can be estimated, and therefore, two observables are missing. One such observable is the microlensing parallax (Refsdal, 1966; Gould, 1992), and another is a measurement of the Einstein radius in angular units. The microlensing parallax can be measured for any event with sufficiently long time scales (e.g., Alcock et al. 1995) or simultaneously using long-baseline space-based observations (e.g., Dong et al. 2007; Zhu et al. 2015). Measuring the Einstein radius in angular units is, however, more difficult. One solution is to use the finite-source effect (Gould, 1994; Witt & Mao, 1994; Nemiroff & Wickramasinghe, 1994), but such events are extremely rare ($\mathord{\sim}10^{-3}$) for BH events. Resolving the lensed images is also difficult because the typical Einstein radius rarely exceeds a few milli-arcseconds (mas). However, this was already demonstrated using interferometric techniques and may become an industry in the coming years (e.g., Dong et al. 2019; Cassan et al. 2022; Wu et al. 2024).
@@
Another approach, which is the main driver for our work, is to measure the position of the center-of-light of the lensed images over time. This motion deviates from pure proper motion combined with parallax, enabling the measurement of the Einstein radius in angular units (Hog et al., 1995; Miyamoto & Yoshii, 1995; Walker, 1995; Dominik & Sahu, 2000). This approach was attempted using ground-based AO observations (e.g., Lu et al. 2016) and space-based observations. So far, one isolated black hole has been found using HST and OGLE observations (Sahu et al. 2022; Lam et al. 2022; Lam & Lu 2023; Mróz et al. 2022).
@@
This balance may shift in the future with the advent of the Nancy Grace Roman Space Telescope, which is expected to expand the scope of space-based astrometric microlensing.
[…]
In this paper, we present our method for precise astrometric measurements of seeing-limited ground-based data. We demonstrate this method on data from the KMTNet telescopes (Kim et al. 2016). This method can provide an astrometric precision of about 0.5 mas on a large sample of events. This astrometric precision is likely enough to detect microlensing events with large Einstein radii. In companion papers, we apply this method to a large set of KMTNet data to measure proper motions and constrain the mass of microlensing events.
@@ 2 Theoretical astrometric precision of seeing-limited observations
Before describing our astrometric reduction procedure, it is worthwhile to discuss the expected astrometric precision. In the best case scenario, the precision of astrometric measurements is limited by Poisson noise, where the astrometric noise (per axis) is roughly $\approx\sigma_{\rm PSF}/(S/N)_{\rm meas}$. Here $\sigma_{\rm PSF}$ is the $\sigma$-width of a Gaussian-like point spread function (PSF; i.e., $\sigma_{\rm PSF}\cong FWHM/2.35$), where the FWHM is the Full-Width at Half Maximum, $(S/N)_{\rm meas}$ is the $S/N$ for a measurement process: $S/N=N_{\rm ph}/\sqrt{(N_{\rm ph}+B)}$, $N_{\rm ph}$ is the number of photons from the source and $B$ is the variance of all the background contributions. For example, for a source with a FWHM of approximately $1^{\prime\prime}$ and a signal-to-noise ratio ($S/N$) of 1000, the Poisson-limited precision predicts an astrometric accuracy of 0.4 mas. However, in reality, Earth’s turbulent atmosphere causes changes in the photon’s direction, and the spatial distribution of photons from a point source is highly correlated on timescales of tens of milliseconds. Furthermore, the phase scintilations have an angular correlation scale of about ten arcseconds in visible light (i.e., the iso-planatic patch). For instance, in the Palomar Transient Factory (PTF, Law et al. 2009), with $\sigma_{\rm PSF}=1^{\prime\prime}$, $S/N\approx 1000$, and a 60-second exposure, the typical astrometric accuracy relative to GAIA is about 14 mas per axis (Ofek 2019), which is approximately 1.5 orders of magnitude worse than the Poisson-limited prediction. This level of accuracy occurs because Poisson noise dominates only when the number of photons from the source, per atmospheric coherence time (roughly 20 ms), is less than one. In the $V$-band, this happens for sources fainter than a magnitude of approximately $20+5\,\log_{10}{D}$, where $D$ is the telescope diameter in meters.
@@
Furthermore, at some level of astrometric precision, we may expect that systematic errors will kick in. There are many potential reasons for systematics, including inhomogeneity in pixel size, sampling effects, color terms, and more (e.g., Beamer et al. 2015).
@@
However, because observations taken over a long period of time are expected to be uncorrelated, averaging multiple epochs can improve the astrometric precision by $1/\sqrt{N_{\rm ep}}$, where $N_{\rm ep}$ is the number of epochs. This was demonstrated, for example, using data from the Large Array Survey Telescope (Ofek et al., 2023a, b). For some KMTNet microlensing events, $N_{\rm ep}\sim 5000$, and therefore, we may expect measurements with a precision on the order of 0.1 mas. In practice, however, some cases may be limited by systematic errors.
@@ 3 The Data
We use observations from the Korea Microlensing Telescope Network (KMTNet; Kim et al. 2016), a global array of three identical 1.6-meter telescopes located at the Cerro-Tololo Inter-American Observatory (CTIO) in Chile, the South African Astronomical Observatory (SAAO) in South Africa, and the Siding Spring Observatory (SSO) in Australia. Each telescope is equipped with a mosaic of four 9k $\times$ 9k CCDs, yielding a field of view of approximately $4~{}\mathrm{deg}^{2}$ and a pixel scale of $0.4~{}\mathrm{arcsec\,pixel^{-1}}$. The typical atmospheric seeing conditions at these sites are between $0.6^{\prime\prime}$$1.0^{\prime\prime}$ at CTIO, $0.8^{\prime\prime}$$1.4^{\prime\prime}$ at SAAO, and $1.0^{\prime\prime}$$2.0^{\prime\prime}$ at SSO (Kim et al., 2016).
KMTNet’s microlensing survey targets the Galactic bulge, covering a total area of $\sim$97 deg2 divided into overlapping fields observed with a tiered cadence strategy. Central fields are monitored at up to 4 hr-1 by each telescope, with overlapping pairs reaching effective cadences of $\sim$8 hr-1. Outer fields are observed less frequently, with cadences as low as 0.2 hr-1 (Kim et al., 2018).
[…]
In this work, we use I-band images from the KMTNet BLG17K0103 field, obtained with the CTIO observatory, to evaluate the pipeline performance and characterize systematics. Additionally, I-band images from the KMTNet BLG15M0306 field, taken with both the CTIO and SAAO observatories, are used to test the consistency between different sites. These two fields were chosen because they are relatively diluted and benefit from good observational cadence, making them well-suited for astrometric analysis and pipeline development.
@@
Table 1: Summary of the observations over the KMTNet fields used in this work. The number of epochs refers to the total number of observations, with the number of epochs used in the astrometric analysis shown in parentheses.
KMTNet field Total # of epochs # CTIO epochs (used) # SAAO epochs (used) Epochs range
BLG17K0103 18336 7671 (6557) 5201 (– ) Feb-2016 – May-2023
BLG15M0306 18527 7737 (6283) 5313 (4173) Feb-2016 – May-2023

### v2

# Towards sub-milliarcsecond astrometric precision using high-cadence seeing-limited imaging

@@ Abstract
The Earth’s atmospheric turbulence degrades the precision of ground-based astrometry. Here we discuss these limitations and propose that, with proper treatment of systematics and by leveraging the many epochs available from the Korean Microlensing Telescope Network (KMTNet), seeing-limited observations can reach sub-milliarcsecond precision. Such observations may be instrumental for the detection of Galactic black holes via microlensing. We present our methodology and pipeline for precise astrometric measurements using seeing-limited observations. The method is a variant of Gaia’s Astrometric Global Iterative Solution (AGIS) that include several detrending steps. Tests on 6,500 images of the same field, obtained by KMTNet with typical seeing condition of 1 arcsecond and pixel scale of 0.4 arcsecond, suggest that we can achieve, at the bright end (mag $\lesssim$17), per-epoch relative astrometric precision of $\mathord{\sim}$5 mas and relative proper motion precision of 0.1-0.2 mas yr-1 over a baseline of approximately five years, using data from the Cerro Tololo Inter-American Observatory (CTIO) site. Time binning on 5–20 day cadences improves the bright-source precision to $\sim$2 mas per coordinate on astrometric microlensing-relevant timescales. The precision is estimated using bootstrap simulations and further validated by comparing results from two independent KMTNet telescopes.
keywords:
astrometry — gravitational lensing: micro — techniques: image processing — stars: black holes — atmospheric effects — proper motions
@@ 1 Introduction
For the past two millennia, precision astrometry provided a plethora of discoveries. However, due to the Earth’s turbulent atmosphere, ground-based astrometry is considerably less efficient (and even less cost-effective) compared with space-based astrometry. The main problems of ground-based astrometry are: (i) The starting point of the precision (i.e., given a single photon) is the seeing rather than the diffraction limit; (ii) Due to atmospheric scintillations, the photons arriving from the stars are spatially correlated, and therefore the measurement precision decrease slower than $1/\sqrt{N_{\rm ph}}$, where $N_{\rm ph}$ is the number of photons. Although adaptive optics observations and interferometric techniques (e.g., genzel1997nature; Ghez+2005ApJ_StarOrbit_BH_GalCenter; gravity_interferometry_2010; dong2019_microlens_resolve_first) can partially solve these problems, they are currently expensive and limited to a small field of view. These two facts make it currently difficult for ground-based astrometry to compete with space-based astrometry. However, even the leading space astrometry mission, Gaia (GAIA+2016_GAIA_mission), has some limitations. Two such limitations are the capability of providing astrometric measurements over short time scales ($\lesssim$ a few months), and the relative inefficiency of Gaia in crowded fields, such as the Galactic bulge. Indeed, several science cases can benefit from precision astrometry measured on time scales shorter than a few months, and/or targets in the Galactic bulge. Examples include searching for astrometric microlensing due to compact objects (e.g., Lu+2016_astrometricMicrolensing_KeckAO; Sahu+2017_astrometricMicrolensing_WD_mass_HST; rybicki_bh_ml_detection_2018; sahu2022isolated), searching for lensed quasars and measuring their time delays (e.g., Springer+Ofek2021_TimeDelayI_FluxOnly; Springer+Ofek2021_TimeDelayII_FluxAstrometry), and searching for binary asteroids (e.g., Segev+2023MNRAS_AstrometricDetection_BinaryAsteroids_GAIADR2).
@@
In this work, we develop and test a method for precision astrometry using high-cadence, ground-based observations of microlensing events towards the Galactic bulge. Our main motivation for the development of this methodology is to find more examples of isolated stellar-mass black holes (BH), and eventually to measure their mass function. This is of great importance for our understanding of the core-collapse supernova explosion mechanism (e.g., Heger+2003ApJ_HowMassiveStarsEndLife_Supernovae_CoreCollapse) and the birth of BH binaries and gravitational wave events (e.g., Abbott+2016_LIGO_FirstGW_BH_Merger). To date, most of the known stellar mass black holes are detected only in binary systems such as X-ray binaries (e.g., BH_xray_binary_mcclintock2003black and xray_binary_ozel2010black) and BH-BH (or BH-neutron star) mergers by Gravitational Waves (e.g., BH_merger_abbott2016observation). Three quiescent stellar-mass black holes were identified through their relatively wide-orbit stellar companions in Gaia observations (elbadri_2023sun_bh_a; elbadri2023red_bh_b; panuzzo2024discovery_bh). To study the abundance of black holes in the Galaxy using binary systems, a complete understanding of the system’s evolution is required. Unfortunately, the evolution of such binary systems is complicated and likely includes physical processes that are not fully understood. Even with a complete understanding, binary-based methods would remain inherently limited, as they trace only those black holes that remain in observable binaries.
@@
Blaes+Madau1993ApJ_AccretingISM_IsolatedNS_Xray and Agol+2002_Xrays_IsolatedBlackHoles suggested that isolated compact objects can be detected via X-ray emission that takes place due to Bondi-Hoyle-like accretion from the interstellar medium (ISM). However, searches for such objects have not revealed good candidates so far (Maoz+1997MNRAS_NewClass_EUV; chisholm2003stellar_no_detection_xray_bh and mereghetti2022xray_followup_bh_ogle_110462). Furthermore, even if such objects are identified, measuring their numbers and masses may prove extremely challenging due to the complexities of accretion physics.
@@
The study of microlensing events is currently the most promising path for detecting isolated stellar-mass black holes. To date, tens of thousands of gravitational microlensing events have been detected by the flux magnification using ground-based telescopes, which observe the Galactic Bulge at high cadence, e.g., OGLE (ogle_mission_udalski1992optical); MOA (MOA_II_abe2008moa); KMTNet (kim2016kmtnet). However, the microlensing light curve only rarely provides enough information for the lens mass to be inferred. There are two approaches to deal with this problem. The first is the statistical approach in which, based on the statistical properties of the microlensing population (e.g., the velocity distribution of lenses and sources) one can infer the mass distribution of a sample of microlensing events.(e.g., Gould2000_MeasuringRemnantMassFunction_Microlensing; mroz_wyrzykowski_2021measuring; mroz2021measuring; rybicki2024analysis). The second approach is breaking the degeneracy and measuring the lens mass. Because, in the simplest case, there are two observables (the magnification and time scale), and five unknowns (lens mass, source distance, lens distance, impact parameter, and relative tangential velocity between the source and the lens), the problem is degenerate. For a Galactic microlensing event, the source distance can be estimated, and therefore, two observables are missing. One such observable is the microlensing parallax (refsdal1966possibility_parallax_satelite; ml_plx_gould1992extending), and another is a measurement of the Einstein radius in angular units. The microlensing parallax can be measured for any event with sufficiently long time scales (e.g., alcock1995first_first_microlens_parallax) or simultaneously using long-baseline space-based observations (e.g., dong2007first_microlens_parallax; zhu2015spitzer_microlens_parallax_mass). Measuring the Einstein radius in angular units is, however, more difficult. One solution is to use the finite-source effect (gould1994proper_finite_source_fs; witt_mao_1994_fs_effect; nemiroff1994_finite_source_fs), but such events are extremely rare ($\mathord{\sim}10^{-3}$) for BH events. Resolving the lensed images is also difficult because the typical Einstein radius rarely exceeds a few milli-arcseconds (mas). However, this was already demonstrated using interferometric techniques and may become an industry in the coming years (e.g., dong2019_microlens_resolve_first; cassan2022microlensing_mass_rotating_resolution; Wu+2024ApJ_Microlensing_Gracity_Resolved).
@@
Another approach, which is the main driver for our work, is to measure the position of the center-of-light of the lensed images over time. This motion deviates from pure proper motion combined with parallax, enabling the measurement of the Einstein radius in angular units (hog1995macho; miyamoto1995astrometry; walker1995microlensed; dominik2000astrometric). Astrometric microlensing is observed as a relative effect, measured through the displacement of the source centroid with respect to the surrounding reference stars. Consequently, our analysis is formulated within a relative astrometric framework. This approach was attempted using ground-based AO observations (e.g., Lu+2016_astrometricMicrolensing_KeckAO) and space-based observations. So far, one isolated black hole has been found using HST and OGLE observations (sahu2022isolated; lam2022isolated_ns_or_bh; lam2023reanalysis; mroz2022systematic).
@@
This balance may shift with the Nancy Grace Roman Space Telescope (Roman), whose Bulge survey will deliver $\mathord{\sim}$15 min cadence across six $\mathord{\sim}$70 day seasons with per-epoch astrometry $\lesssim$1 mas. Yet months-long gaps between seasons leave mid-duration ($\mathord{\sim}$60–70 day) events under-sampled; ground surveys can supply dense coverage and early/late baselines. In that sense, our survey remains competitive in discovery volume and temporal coverage, and strongly complementary for characterization: it extends time baselines across Roman seasons and provides an independent consistency check on Roman astrometry. Because our program precedes Roman, it is both a test bed and a complement.
To date, microlensing surveys (e.g., ogle_mission_udalski1992optical; MOA_II_abe2008moa; kim2016kmtnet) detect $\mathord{\sim}$3000 verified microlensing events per year. Gould2000_MeasuringRemnantMassFunction_Microlensing estimates that $\mathord{\sim}$1% of all microlensing events are due to stellar-mass black holes. Therefore, we expect that microlensing samples include numerous events in which the lens is an isolated BH, and the direct measurement of the Einstein radius may allow us to search for these events. However, the detection of the astrometric microlensing signal requires typical precision which are of the order of milliarcseconds or below. As far as we know, such precisions using seeing-limited ground-based astrometry were not reported in the past, although several attempts came close to the level of a few milliarcsecond precision (e.g., monet1983ccd; Soszynski+2002AcA_OGLE_ProperMotion_LMC_SMC; Sumi+2004MNRAS_OGLE_ProperMotion_GalacticBuldge).
In this paper, we present our method for precise astrometric measurements of seeing-limited ground-based data. We demonstrate this method on data from the KMTNet telescopes (kim2016kmtnet). This method can provide an astrometric precision of about 0.5 mas on a large sample of events. This astrometric precision is likely enough to detect microlensing events with large Einstein radii. In companion papers, we apply this method to a large set of KMTNet data to measure proper motions and constrain the mass of microlensing events.
The astrometric microlensing signal typically extends over several times the Einstein crossing time $t_{\mathrm{E}}$, i.e., months to years depending on lens mass and relative kinematics. While the photometric magnification is confined to a period of order $t_{\mathrm{E}}$, the centroid shift remains detectable well before and after peak. For our purposes, the critical requirement is therefore not decade-long baselines, but sufficiently dense sampling across the times of maximum deflections as well as observations obtained a few Einstein crossing times before and after the maximum magnification, in order to constrain the proper motion. The KMTNet survey provides exceptionally high cadence in the Galactic bulge ($\mathord{\sim}$10–200 epochs per field per day), enabling us to combine hundreds of measurements into effective astrometric epochs; in the ideal limit of uncorrelated noise, the centroid precision scales as $1/\sqrt{N}$. For example, binning 500 exposures (per-exposure precision $\sim$5 mas) over a few days yields a nominal precision of $\approx 5/\sqrt{500}\simeq 0.22$ mas. In practice, residual systematics and short-timescale correlations limit this gain ($N\rightarrow N_{\rm eff}\leq N$). Even so, over a typical months-long event, this strategy yields the equivalent of hundreds of high-precision astrometric epochs, hopefully sufficient to measure $\theta_{\mathrm{E}}$ with the accuracy required for our science goals.
@@ 2 Theoretical astrometric precision of seeing-limited observations
Before describing our astrometric reduction procedure, it is worthwhile to discuss the expected astrometric precision. In the best-case scenario, the precision of astrometric measurements is bounded by the Cramér–Rao inequality; in the photon-limited regime, this bound reduces to the familiar Poisson limit, where the astrometric noise (per axis) is roughly $\approx\sigma_{\rm PSF}/(S/N)_{\rm meas}$. Here $\sigma_{\rm PSF}$ is the $\sigma$-width of a Gaussian-like point spread function (PSF; i.e., $\sigma_{\rm PSF}\cong FWHM/2.35$), where the FWHM is the Full-Width at Half Maximum, $(S/N)_{\rm meas}$ is the $S/N$ for a measurement process: $S/N=N_{\rm ph}/\sqrt{(N_{\rm ph}+B)}$, $N_{\rm ph}$ is the number of photons from the source and $B$ is the variance of all the background contributions. For example, for a source with a FWHM of approximately $1^{\prime\prime}$ and a signal-to-noise ratio ($S/N$) of 1000, the Poisson-limited precision predicts an astrometric accuracy of 0.4 mas. However, in reality, Earth’s turbulent atmosphere causes changes in the photon’s direction, and the spatial distribution of photons from a point source is highly correlated on timescales of tens of milliseconds. Furthermore, the phase scintilations have an angular correlation scale of about ten arcseconds in visible light (i.e., the iso-planatic patch). For instance, in the Palomar Transient Factory (PTF, PTF_paper_law2009palomar), with $\sigma_{\rm PSF}=1^{\prime\prime}$, $S/N\approx 1000$, and a 60-second exposure, the typical astrometric accuracy relative to GAIA is about 14 mas per axis (Ofek2019_Astrometry_Code), which is approximately 1.5 orders of magnitude worse than the Poisson-limited prediction. This level of accuracy occurs because Poisson noise dominates only when the number of photons from the source, per atmospheric coherence time (roughly 20 ms), is less than one. In the $V$-band, this happens for sources fainter than a magnitude of approximately $20+5\,\log_{10}{D}$, where $D$ is the telescope diameter in meters.
State-of-the-art space telescopes, e.g., HST and JWST, deliver sub-milliarcsecond astrometry. For HST, bulge microlensing analyses achieve $\sim$0.2-0.3 mas precision per epoch sahu2022isolated; lam2022isolated_ns_or_bh; for JWST, the NIRCam geometric-distortion solution has precision at the $\lesssim$0.2 mas level (griggio2023photometry).
@@
Furthermore, at some level of astrometric precision, we may expect that systematic errors will kick in. There are many potential reasons for systematics, including inhomogeneity in pixel size, sampling effects, color terms, and more (e.g., LSST_tree_rings_beamer2015study).
@@
However, because observations taken over a long period of time are expected to be uncorrelated, averaging multiple epochs can improve the astrometric precision by $1/\sqrt{N_{\rm ep}}$, where $N_{\rm ep}$ is the number of epochs. This was demonstrated, for example, using data from the Large Array Survey Telescope (Ofek+2023PASP_LAST_Overview; Ofek+2023PASP_LAST_PipeplineI). For some KMTNet microlensing events, $N_{ep}\mathord{\sim}5000$, comparable to the number of epochs in the fields analyzed in this work, and therefore we may expect measurements with a precision on the order of 0.1 mas. Since most KMTNet events are monitored in denser fields with $N_{ep}\mathord{\sim}15000$, the achievable precision in those cases could be even better. In practice, however, some cases may be limited by systematic errors.
@@ 3 The Data
We use observations from the Korea Microlensing Telescope Network (KMTNet; kim2016kmtnet), a global array of three identical 1.6-meter telescopes located at the Cerro-Tololo Inter-American Observatory (CTIO) in Chile, the South African Astronomical Observatory (SAAO) in South Africa, and the Siding Spring Observatory (SSO) in Australia. Each telescope is equipped with a mosaic of four 9k $\times$ 9k CCDs, yielding a field of view of approximately $4~\mathrm{deg}^{2}$ and a pixel scale of $0.4~\mathrm{arcsec\,pixel^{-1}}$. The typical atmospheric seeing conditions at these sites are between $0.6^{\prime\prime}$$1.0^{\prime\prime}$ at CTIO, $0.8^{\prime\prime}$$1.4^{\prime\prime}$ at SAAO, and $1.0^{\prime\prime}$$2.0^{\prime\prime}$ at SSO (kim2016kmtnet).
KMTNet’s microlensing survey targets the Galactic bulge, covering a total area of $\sim$97 deg2 divided into overlapping fields observed with a tiered cadence strategy. Central fields are monitored at up to 4 hr-1 by each telescope, with overlapping pairs reaching effective cadences of $\sim$8 hr-1. Outer fields are observed less frequently, with cadences as low as 0.2 hr-1 (kim2018kmtnet).
KMTNet observes in both I-band (most of the time) and V-band (every $\mathord{\sim}$10 epochs). Over the eight years of 2016–2023, each field has accumulated between a few thousand and over $10^{5}$ images, depending on its assigned cadence. The network collects approximately 50,000–75,000 science images annually during the bulge season.
In this work, we use I-band images from the KMTNet BLG17K0103 field, obtained with the CTIO observatory, to evaluate the pipeline performance and characterize systematics. Additionally, I-band images from the KMTNet BLG15M0306 field, taken with both the CTIO and SAAO observatories, are used to test the consistency between different sites. These two fields were chosen because they are relatively diluted and provide a cadence that is well-suited for astrometric analysis and pipeline development. Over the 2016–2023 baseline, BLG17K0103 contains about 7,000 images from CTIO, while the denser microlensing fields typically accumulate closer to 15,000 images. However, despite their higher cadence, the more crowded fields are less favourable for precise astrometric analysis and calibration of the pipeline, due to blending and stronger systematics.
@@
Table 1: Summary of the observations over the KMTNet fields used in this work. The number of epochs refers to the total number of observations, with the number of epochs used in the astrometric analysis shown in parentheses (after filtering for airmass and removal of poor-quality images).
KMTNet field Total # of epochs # CTIO epochs (used) # SAAO epochs (used) Epochs range RA [hms] Dec [dms]
BLG17K0103 18336 7671 (6557) 5201 (– ) Feb-2016 – May-2023 17:43:56.54 -32:52:15.42
BLG15M0306 18527 7737 (6283) 5313 (4173) Feb-2016 – May-2023 17:40:04.17 -26:02:36.60